OpenStudy (anonymous):

Find the slope and y intercept of 8x=2y+7

hartnn (hartnn):

can you isolate y from 8x=2y+7 ?

hartnn (hartnn):

like bring it in the form of y=mx+c

OpenStudy (anonymous):

Like -2y=7-8x

hartnn (hartnn):

yes, and now divide both sides by -2

OpenStudy (anonymous):

So then my slope is -3.5 and y intercept is 4 correct?

hartnn (hartnn):

y= 4x-3.5

hartnn (hartnn):

in y=mx+c m=slope, so here slope = 4

OpenStudy (anonymous):

So then my slope is 4 and y intercept is -3.5 and I had it backwards before

hartnn (hartnn):

yes, correct.

OpenStudy (anonymous):

ok thank you

hartnn (hartnn):

welcome ^_^